The perimeter of a square is three times the perimeter of a regular hexagon. The length of a side of the square is 2 inches more than four times the length of a side of the hexagon. What is the perimeter of the square and hexagon?
Answer by scott8148(6628)

let x="perimeter of square", and y="perimeter of hexagon"

"The perimeter of a square is three times the perimeter of a regular hexagon" __ x=3y

"The length of a side of the square is 2 inches more than four times the length of a side of the hexagon"
__ x/4=2+4(y/6)

substituting __ 3y/4=2+4(y/6)

multiplying by 12 __ 9y=24+8y